OpenStudy (anonymous):

Simplify using properties of logarithms as a single logarithm. 1/3log7(8)

OpenStudy (anonymous):

Can you please clarify the problem?

OpenStudy (anonymous):

Simplify using properties of logarithms to write each expression as a single logarithm. Do not evaluate using a calculator. (b) 1/3log7(8) (c) 2log(x) + 3log(x+1) (d) log(x) - 1/2log(y)

OpenStudy (anonymous):

No I mean I have no clue how it's written, is that log base 7?

OpenStudy (anonymous):

\[\huge \frac{ 1 }{ 3 }\log _{7}8?\]

OpenStudy (anonymous):

Like that?

OpenStudy (anonymous):

I'm not sure how to use the equation editor... so I try to write it out by hand but usually mess it up

OpenStudy (anonymous):

yes, that's it

OpenStudy (anonymous):

\[n*\log _{b}(m) =>\log _{b}(m^n)\]

OpenStudy (anonymous):

\[\log_{7}(8)^{1/3}\]

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

the 1/3 should be inside the ()

OpenStudy (anonymous):

Yeah so what's 8^(1/3)?

OpenStudy (anonymous):

You have a fraction in the exponent, so it would be \[\sqrt[3]{8}\]

OpenStudy (anonymous):

I think

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Yeah \[\frac{ \log(2) }{ \log(7) }\]

OpenStudy (anonymous):

how do you solve that without a calc?

OpenStudy (anonymous):

You leave it as that

OpenStudy (anonymous):

oh, that makes a lot of sense